Identify the t-value whose number of samples n = 25, area = 0.01




n = 30 area = 0.005




n = 27 , area = 0.0025




n = 35 area = 0.0010




USE A SHORT BOND PAPER




n = 40 area = 0.0005




SHOW YOUR SOLUTIONS

1
Expert's answer
2022-05-19T17:00:05-0400

a.


"df=n-1=24,"

"t=-2.492159"

"P(t<-2.492159)=0.05"

b.


"df=n-1=29,"

"t=-2.756386"

"P(t<-2.756386)=0.005"

c.


"df=n-1=26,"

"t=-3.066909"

"P(t<-3.066909)=0.0025"

d.


"df=n-1=34,"

"t=-3.347934"

"P(t<-3.347934)=0.0010"

e.


"df=n-1=39,"

"t=-3.55812"

"P(t<-3.55812)=0.0005"